OpenStudy (anonymous):

Write each equation in slope - intercept form. Then identify the slope and the y - intercept for each line. 2x - y = 12

OpenStudy (anonymous):

2x -y = 12 => -2x +y = -12 => y = 2x -12 slope = 2, intercept = -12

OpenStudy (anonymous):

ty